# Voltage-sensitive distribution factors for contingency analysis and topology optimization

### v1

@@ V-C Example
We demonstrate the application of the generalized bus split distribution factors. We start from an AC load-flow of the base case to obtain the reference values $\hat{v}_{m}$ and $\hat{\theta}_{m}$ and the matrices ${\boldsymbol{M}}_{c}$ and its inverse ${\boldsymbol{M}}_{c}^{-1}$. The impact of all topology changes can be readily computed via low rank updates: Finite topology changes can be treated via Eq. (10), the merging of buses and closing of switches via Eq. (17) and bus splits via Eq. (20).
@@
Here, we explore the numerical results and accuracy of the generalized bus split distribution factors for an elementary test (Fig. 4). We simulate the opening of a busbar coupler at substation VLevel3 comparing the generalized DC+ approximation given by Eq. (20) to full AC load-flow computation. The split results in a voltage drop at bus bar (BB1) from $225\text{\,}\mathrm{k}\mathrm{V}$ to $198\text{\,}\mathrm{k}\mathrm{V}$ and a phase angle drop from $0.2\text{\,}{}^{\circ}$ to $14\text{\,}{}^{\circ}$. These values are accurately reproduced by the generalized DC+ approximation that predicts a voltage magnitude drop to $201\text{\,}\mathrm{k}\mathrm{V}$ and a voltage angle drop to $12\text{\,}{}^{\circ}$.
VI Conclusion and Outlook
@@
We have introduced a framework for voltage-sensitive distribution factors based on a linearized AC load-flow formulation. This approach enables efficient evaluation of contingency scenarios and dynamic topology modifications, including line outages and busbar operations. The DC+ framework enables further extensions, such as the modeling of Static Synchronous Compensators (STATCOMs), as the formulation of the $M$ matrix aligns well with fixed Jacobian approaches.
@@
The current implementation demonstrates the feasibility of approximating voltage and reactive power responses with good accuracy, while providing more accurate active power and voltage angle responses than the traditional DC approximation. While the linearization introduces limitations, particularly in reactive power estimation and in scenarios with large deviations from the reference state, the results show strong alignment with full AC simulations across small and large test grids.
The DC+ framework is based on low-rank updates via the Woodbury matrix identity, making it well-suited for GPU acceleration [35]. Hence, it is a promising candidate for all time-critical applications such as heuristic algorithms for grid topology optimization.

### v2

@@ V-C Example
We demonstrate the application of the generalized bus split distribution factors. We start from an AC power flow of the base case to obtain the reference values $\hat{v}_{m}$ and $\hat{\theta}_{m}$ and the matrices ${\boldsymbol{M}}_{c}$ and its inverse ${\boldsymbol{M}}_{c}^{-1}$. The impact of all topology changes can be readily computed via low-rank updates: Finite topology changes can be treated via Eq. (10), the merging of buses and closing of switches via Eq. (17) and bus splits via Eq. (20).
@@
Here, we explore the numerical results and accuracy of the generalized bus split distribution factors for an elementary test (Fig. 4). We simulate the opening of a busbar coupler at substation VLevel3 comparing the generalized DC+ approximation given by Eq. (20) to full AC power flow computation. The split results in a voltage drop at bus bar (BB1) from $225\text{\,}\mathrm{k}\mathrm{V}$ to $198\text{\,}\mathrm{k}\mathrm{V}$ and a phase angle drop from $0.2\text{\,}{}^{\circ}$ to $14\text{\,}{}^{\circ}$. These values are accurately reproduced by the generalized DC+ approximation that predicts a voltage magnitude drop to $201\text{\,}\mathrm{k}\mathrm{V}$ and a voltage angle drop to $12\text{\,}{}^{\circ}$.
VI Numerical Performance
The generalized distribution factors enable new applications in power system operation and planning that require extraordinarily high computational throughput. For example, topology optimization using metaheuristics such as genetic algorithms may involve evaluating hundreds of thousands of switching configurations. When each candidate must additionally undergo a full N1N-1 contingency screening, the resulting workload easily grows to hundreds of millions of power flow evaluations. Similar requirements arise in transmission expansion planning, where many corridor options and operating scenarios must be assessed under large contingency sets, and in operational settings where operators may need to rapidly evaluate large libraries of remedial actions. Importantly, these workloads are inherently parallelizable because each topology and contingency can be evaluated independently, making them well suited for modern GPU architectures.
Meeting these demands is out of reach for nonlinear AC power flow based approaches, whose computational cost is typically orders of magnitude too high for such outer-loop search procedures. The classical DC approximation is computationally much faster and can be implemented very efficiently on GPUs, enabling extremely high evaluation rates [35]. However, it cannot capture voltage magnitude variations and reactive power effects and therefore becomes unreliable when voltage security and stability constraints become binding. The generalized distribution factors introduced in this work aim to bridge this gap, combining much of the computational efficiency of linear methods with a significantly improved representation of voltage-sensitive phenomena.
We benchmark the numerical performance of the generalized distribution factors as follows. For various standard test grids we compute the state variables in the N0N-0 case using a standard Newton-Raphson solver. Then we do an N1N-1 contingency analysis using the methods introduced in Sec. IV-A. That is, we compute the state variables $\theta_{n}$ and $v_{n}$ via Eq. (9) as well as the currents for all possible non-islanding single branch outages and record the wall time. This is repeated several times to minimize statistical errors. As a figure of merit, we report the number of power flows per second which is defined as
$\mathcal{N}=\frac{n_{\mathrm{outages}}\times n_{\mathrm{repetitions}}\times\mathrm{batch\_size}}{\mathrm{walltime}}\,.$ (21)
We set $n_{\mathrm{repetitions}}=500$ for all GPU experiments and $n_{\mathrm{repetitions}}=3$ for all CPU experiments. Parameters of the test grids are summarized in Tab. I.
TABLE I: Benchmark Summary. Properties of the test systems and numerical performance for optimized batch size. Performance is computed according to Eq. (21) and reported in power flows per second.
Test System GBnetwork case2869 case1354 case118
$n_{\text{buses}}$ 2224 2869 1354 118
$n_{\text{outages}}$ 2083 3579 1288 177
$n_{\text{branches}}$ 3207 4582 1991 186
Generalized distribution factors (proposed method)
Load time [s] 4.72 6.84 2.56 0.92
Power flows/s $3.1\!\times\!10^{6}$ $1.5\!\times\!10^{6}$ $3.9\!\times\!10^{6}$ $38.7\!\times\!10^{6}$
Optimal batch size 1 1 128 1024
Newton–Raphson (lightsim2grid)
Power flows/s (1 iter) 308 192 525 8688
Power flows/s (20 iter) 98 66 176 3764
We use PyPowSyBl [33] and pandapower [36] for grid file loading and the initial Newton-Raphson process plus some preprocessing routines in native Python. The high-performance N1N-1 loop using the generalized distribution factors is implemented in JAX [37]. All tests were run on a virtual machine with 48 AMD EPYC 7V13 cores, 440 GB RAM and one NVIDIA A100 GPU with 80GB VRAM. Correctness is verified by comparing the results to a single-iteration Newton–Raphson step with the initial guess of the N0N-0 result. Our comparison baseline is lightsim2grid [38] on a single CPU core. We run either one or 20 Newton-Raphson iterations with an initial guess from the N0N-0 computation. We emphasize that running one iteration is comparable in accuracy to the generalized distribution factors introduced in this article.
The numerical results demonstrate that the generalized distribution factors lend themselves well to massively parallel execution on modern GPU architectures. Even for large transmission networks with several thousand buses, our implementation achieves more than one million power flow evaluations per second, while smaller systems reach several tens of millions of evaluations per second. Compared to the AC-based benchmark lightsim2grid, the proposed method improves numerical throughput by several orders of magnitude. When compared to a single Newton–Raphson iteration, which yields an approximation of comparable accuracy, the speedup reaches roughly four orders of magnitude, primarily due to the massive parallelism enabled by GPU execution. Compared to a full Newton–Raphson solve with 20 iterations, corresponding to a fully converged AC power flow solution, the performance gap becomes even larger. The approach comes at the cost of several seconds of preprocessing for matrix construction and grid loading before the first results can be generated, making it unsuitable for workloads with fewer than a few hundred cases. This substantial speedup highlights the advantage of linearized distribution-factor approaches for large-scale contingency screening and topology analysis.
As expected, the computational performance is lower than that of the plain DC formulation due to the additional state variables and the larger Jacobian matrix involved. In earlier work on GPU-accelerated DC distribution factors, we reported throughputs of up to billions of power flows per second under comparable workloads [35]. Nevertheless, the present method retains much of the computational efficiency of DC-based techniques while providing a substantially improved representation of voltage-sensitive phenomena.
Overall, these results show that the proposed generalized distribution factor formulation, combined with efficient GPU parallelization, enables extremely fast screening of contingencies and topology modifications. This capability opens the door to new applications in topology optimization and transmission planning, where large numbers of candidate topologies must be evaluated under extensive contingency sets.
VII Conclusion and Outlook
@@
We have introduced a framework for voltage-sensitive distribution factors based on a linearized AC power flow formulation. This approach enables efficient evaluation of contingency scenarios and dynamic topology modifications, including line outages and busbar operations. The DC+ framework enables further extensions as the formulation of the matrix $\boldsymbol{M}$} aligns well with fixed Jacobian approaches.
@@
The current implementation demonstrates the feasibility of approximating voltage and reactive power responses with good accuracy, while providing more accurate active power and voltage angle responses than the traditional DC approximation. While the linearization introduces limitations, particularly in reactive power estimation, the results show strong alignment with full AC simulations across different test grids. Further studies suggest that deviations are largest when the voltages violate operational limits. Even in these cases, DC+ provides valuable insights as it typically predicts the occurrence of violations while it cannot capture the precise magnitude.
The DC+ framework is based on low-rank updates via the Woodbury matrix identity, making it well-suited for GPU acceleration [35]. Hence, it is a promising candidate for time-critical applications such as heuristic algorithms for grid topology optimization. In our GPU implementation we achieve more than one million power flow evaluations per second even for large transmission networks with several thousand buses. These results demonstrate that voltage-sensitive contingency screening and topology modification analysis can be performed at extremely high throughput, enabling large-scale search and optimization procedures that would be computationally prohibitive with nonlinear AC power flow methods.